# Understanding Ethical Practices in AI: Insights from a Cross-Role, Cross-Region Survey of AI Development Teams

### v1

@@ 5.1. Demographics
After filtering, we accepted 414 responses from our survey (see Section 3.1.2). 201 participants worked at companies in North America, 118 in EU+EEA+UK, and 95 in other or non-specified locations. $\sim$74% of them indicated they had experience developing AI products (i.e., Group B) while $\sim$26% did not have any experience (i.e., Group A). Table 1 illustrates that the majority are in AD (e.g., AI/software developer) (47.1%) and AM (e.g., AI manager) (18.4%) roles. Table 2 shows that most participants identify as male (67.6%), are below the age of 45 ($\sim$88%), and have completed a Bachelor’s or higher degrees ($\sim$89%), with $\sim$77% in computer science, information science, software engineering, data science, and electrical, computer, privacy, or security engineering. Among the 13.1% with a business degree, $\sim$32% are in AD, $\sim$30% in AM, and $\sim$24% in RA roles. Of the 10.2% who did not have degrees within any of the above groups, $\sim$25% were in health sciences and $\sim$10% in law and psychology (among others). Most participants ($\sim$60%) have at least two years of experience in AI development.
Table 2. Breakdown of Participants’ Demographics
Age 18-25 (13.0%) 26-35 (50.5%) 36-45 (24.4%) 46+ (11.8%) PnS (0.3%)
5.2. AI Perceptions and Usage
@@
To address RQ1, we examine participants’ perceptions of AI, its applications, and its impact on their work.
5.2.1. Defining AI
@@
To understand how participants perceive AI, we asked all of them to define AI within their industry. We used the NIST Taxonomy of Human-AI Use Cases (Theofanos et al., 2023) to categorize their responses (see Appendix A - Table 22 for more details on the taxonomy and Figure 1 for a breakdown of participants’ responses). This taxonomy does not capture some technical definitions, such as machine learning. We subsequently created a dedicated category for Machine/Deep Learning (ML/DL).
@@
Figure 1. AI Definitions Based on NIST Human-AI Taxonomy
@@
A significant number of participants in both groups (12.26% and 17.70% for Groups A and B, respectively) highlighted a wider range of use cases, which we could not map to any of the NIST categories. For instance, one participant defined AI as “It is revolutionary and can change everything in this industry.” We categorized these definitions as “Others.” Specifically, in Group B, some of these responses focus on societal or future impacts of AI (e.g., “corruptive and hugely beneficial in equal measures”). Our results indicate that those with experience in AI development (i.e., Group B) identify more technical aspects of AI and its abilities in areas such as data analysis and automation (e.g., “It is an indispensable tool for fast and efficient work that analyzes and verifies a huge volume of data. It performs work that would otherwise be impossible”).
@@
We further analyzed Group B participants based on their roles. “Process automation” is the most common AI definition among AM roles (41.03%), but it is considerably less common among ISec (20%) and AR (18.18%) roles. This may be because the management team is more heavily involved in process improvement, an area where AI-driven automation can have a positive impact. Participants in QA and AD roles are most likely to describe AI as “content synthesis,” which is a more technical definition. QAs are also considerably more likely than other roles to mention “decision making” (27.78%). ISec roles are more than three times more likely to say “content creation” (33.33%) than other roles. This may suggest that individuals in these roles are less likely to be involved in AI development activities, and like Group A participants, see AI through the lens of contemporary narratives rather than its technical aspects. Consistent with prior work, these findings underscore the need for tools and educational frameworks focused on privacy and security aspects of AI, as ISec team members play a critical role in addressing emerging threats that developers are often unaware of (Gao et al., 2024; Meisenbacher et al., 2024).
5.2.2. AI Usage
@@
We asked both Groups A and B about the primary applications of AI they envision in their products. As shown in Table 3, “chatbots, personal assistants, or recommender systems” are considered the most common AI applications across both groups. Group B considers “robotics and automation,” “cybersecurity,” “computer vision,” “human resources,” and “legal” more than Group A. In comparison, Group A participants are $\sim$13%, $\sim$13%, and $\sim$10% more likely to choose “translation and text generation,” “healthcare,” and “chatbots, personal assistants, or recommender systems.” These results further indicate that those without AI development experience (Group A) are more optimistic about applications tailored towards the general public, such as chatbots, while those experienced in developing AI (Group B) consider more job-related or technical applications.
Table 3. Applications of AI in Products
@@
Group B participants could select multiple AI tools and technologies they use in their products, as shown in Figure 2. “Google Cloud AI” and “Model Provider APIs, such as OpenAI, Anthropic, LLama, etc.” are the most popular AI tools selected by the participants (49.50% and 42.52%, respectively). This indicates that development teams primarily use pre-trained models instead of building or training their own. In addition, “Azure ML,” “Pytorch,” “AWS Sagemaker,” and “Tensorflow” are more popular than “transformer-based LLMs” and other deep and machine-learning architectures.
Figure 2. Group B - Tools and Technologies Used in Product or Development.
All participants were asked to discuss how often they use AI tools and what they would see as their main impact on their work. Tables 4 and 5 show the frequency of AI usage and its impact. Most participants (i.e., 82.24% of Group A and 96.67% of Group B) use AI at least 1-3 times per week, while more than 50% of Group B uses AI tools daily. All participants consider “increased efficiency” (71.0% and 77.5% of Groups A and B) and “improved accuracy” (41.1% and 55.3% of Groups A and B) in their work. Interestingly, Group B selected “cost reduction” as the third most crucial impact of AI, with 22.2% more respondents than Group A. In contrast, Group A identified “innovation and creativity” as the third most crucial impact. The results show that all participants recognize AI’s benefits in enhancing workplace tasks, with experienced users more frequently considering automating cost-intensive tasks.
Table 4. Frequency of AI Tool Usage
@@
Table 5. Effects of AI-Enhanced Tools on Development Process and Workflow
@@
Since Group A’s companies did not adopt AI in their development processes, we asked them to select factors that limit their companies’ involvement with AI. As shown in Figure 3, “privacy and security concerns” (45.19%), “lack of trust” (38.46%), and “lack of expertise”(35.58%) are the most significant factors, while “human resources” (16.35%) and “proprietary reasons” (20.19%) appear not as decisive. This indicates the importance of ethical considerations, including privacy, security, and trust, for companies that do not consider AI in their products. As a follow-up question, they had to describe why these factors limit their companies’ involvement. We received a diverse set of 106 responses, in which we identified seven common themes (see Table 6). These responses reinforced the previous finding, with $\sim$29% of participants expressing concerns regarding security and privacy, where many reiterated worries about data sharing and consent (e.g., “customer data concerns”) or were uncertain about the security of certain AI systems (e.g., “if you have proprietary information and materials, you don’t want to feed it into an LLM model that will later regurgitate that information to any user”). 18.87% considered “high costs” as the main factor, while 12.26%, 11.32%, and 10.38%, respectively, mentioned a lack of “experience,” “established AI applications,” or “trust with AI systems.”
Table 6. Group A - Reasons for Limiting AI Usage in Development
@@
Figure 3. Factors Limiting AI Adoption in Development Processes
All participants, regardless of their experience in AI development, primarily define AI in terms of process automation or performance enhancement, report daily use of AI tools, and perceive a positive impact on efficiency and accuracy. Participants with experience in AI development activities tend to engage with more technical aspects of AI in their work, whereas those without such experience typically utilize AI tools, such as chatbots or recommender systems, often for purposes related to innovation and creativity. Concerns regarding security and privacy remain the primary reasons some companies have not yet adopted AI tools in their development processes.
5.3. AI Ethics Familiarity, Knowledge, and Practice
To address RQ2 and RQ3, we assess participants’ familiarity with AI ethics principles (Wörsdörfer, 2025) and governance initiatives, and examine how these are considered across various demographics.
5.3.1. Familiarity with AI Ethics Principles
We asked participants to rate their familiarity with various AI ethics principles, as shown in Figure 4. We combine the responses of somewhat, moderately, and extremely familiar. The results show that both groups are most familiar with “data protection and the right to privacy” (81.2% and 92.3%) and least familiar with “democracy and rule of law” (58.0% and 75.7%). This may be because the concept of privacy is more tangible for development teams. Those with AI development experience are more familiar with “transparency and explainability of AI systems” (89.3%) and “accountability and responsibility” (88.0%), while those in Group A are more familiar with “respect for human rights” (74.8%). For all other principles, Group A’s familiarity is below 70%.
(a) Group A - Familiarity with AI Ethics Principles
(b) Group B - Familiarity with AI Ethics Principles
Figure 4. Degree of Familiarity with AI Ethics Principles
@@ 5.3.1. Familiarity with AI Ethics Principles
Figure 5. Group B - Average Familiarity With AI Ethics Principles by Demographics
To examine whether our analysis of demographics is statistically significant, we conducted a Kruskal-Wallis test on the Likert scale data for the degrees of familiarity. We tested whether there is a correlation between various demographic factors, such as role (H1a), location (H1b), company type (H1c), company size (H1d), and gender (H1e), and the average familiarity with AI ethics principles. As shown in Table 7, we found no evidence of a correlation between demographics and average familiarity among Group A participants. However, without applying the Bonferroni correction, we observe correlations between location, company type, and gender and Group B’s average familiarity (pvalue=0.0102,0.0463p-value=0.0102,0.0463 and $0.003$). When the Bonferroni correction is applied, only the correlation with gender remains statistically significant.
Table 7. P-Values for Hypothesis H1a to H1e
@@
Our analysis of participants’ familiarity with AI ethics principles complements prior work emphasizing the importance of diversity in demographics and expertise within AI development teams (Pant et al., 2024; Vakkuri et al., 2022; Olson et al., 2025). Others have also highlighted the need for effective communication among developer teams (Stahl et al., 2022), and the critical role of management in engaging their teams and stakeholders in ethical decision-making (Kawakami et al., 2024). Involving stakeholders from diverse locations, roles, expertise, genders, and other demographic backgrounds enhances development teams’ capacity to make informed and ethical decisions regarding AI.
5.3.2. Practice and Experience with AI Ethics Principles
@@
In addition to understanding the degrees of familiarity with AI ethics principles (i.e., RQ2), we evaluated if Group B participants consider these principles in the actual AI development lifcycle (i.e., RQ3). The results are shown in Figure 6. As expected, “data protection and the right to privacy” is considered the most important. This is probably because the concept of digital privacy has existed for much longer than other computer-ethical aspects. “Transparency and explainability of AI systems” is regarded as the second most important principle, while “democracy and rule of law” is considered the least important. The results match the degree of familiarity, as seen in Figure 7(b). We tested whether there is a correlation between the average familiarity and the use of principles in AI development (H2a). With the pvalue=0.0000p-value=0.0000, we found that there is a correlation.
@@
In a more role-specific analysis, we noticed that those in AM, AD, AR+AE, and RA roles consider a variety of principles in their AI development activities. For example, those in AM roles consider the principles of “respect for human rights,” “non-discrimination and freedom of privileges,” “harm prevention and beneficence,” and “fairness and justice” whereas those in AR+AE ensure that the “accountability and responsibility” principle is applied in the workflow. In contrast, those in QA or ISec roles do not consider these principles as much as the others, except for “data protection and the right to privacy.” This highlights areas for improvement, as QA roles are crucial for maintaining trust in AI systems (Ala-Luopa et al., 2024) and collaboration with ISec teams is essential to ensure the security and privacy of AI models (Prybylo et al., 2024). Participants working in academic/research institutions tend to prioritize more technical principles, such as “data protection and the right to privacy,” “fairness and justice,” and “transparency and explainability of AI systems,” when developing or researching AI models or datasets. In contrast, those employed in government agencies are likelier to emphasize abstract principles, including “respect for human rights,” “non-discrimination and freedom of privileges,” and “harm prevention and beneficence.” Participants in EU+EEA+UK consider “data protection and the right to privacy” and “democracy and rule of law” more than the participants in North America and other locations. This is expected since privacy is considered a fundamental (human) right and is regulated by the General Data Protection Regulation (GDPR) (European Union, 2024). In line with the above results, female participants consider AI ethics principles in their AI development activities more than their male counterparts.
Figure 6. Group B - Ethics Principles in AI Development Activities
@@
We asked Group B how often they consider AI ethics principles in their work. Most participants consider them “Always” (23.92%) or “Often” (44.52%), while only 9.97% and 2.33%, respectively, consider them “Rarely” or “Never.” Our results show that those in QA (92.3%), ISec (91.7%), AM (91.5%), and AR+AE (90%) roles are more likely to consider a subset of AI ethics principles in their work at least occasionally, whereas those in AD (85.4%) and RA (75%) roles report slightly lower engagement with these principles. Participants with a Bachelor’s degree or higher are $\sim$10% more likely to consider AI ethics principles in their work than those without. Aligned with other observations, female participants include AI ethics principles in their work $\sim$13% more than their male counterparts.
@@
Our results highlight that the degree of familiarity with AI ethics principles - and the extent to which they are considered in practice - varies across demographic factors. Participants in AM roles demonstrate greater familiarity with most principles and are more likely to incorporate them into their work. In contrast, those in QA or ISec roles tend to be the least familiar with these principles, with the exception of “data protection and the right to privacy.” Additionally, both location and company type influence levels of familiarity and application. Overall, female participants not only reported higher familiarity with all principles but also indicated more consistent integration of these principles into their workflows.
5.3.3. Familiarity with AI Governance Initiatives

### v2

@@ 4.1 Demographics
After filtering, we accepted 414 responses (see Section 3.1.2). 201 participants worked at companies in North America, 118 in EU+EEA+UK, and 95 in other or unspecified locations. $\sim$74% of them indicated they had experience developing AI products (i.e., Group B) while $\sim$26% did not have any experience (i.e., Group A). Appendix A - Table 1 illustrates that the majority are in AD (e.g., AI/software developer) (47.1%) and AM (e.g., AI manager) (18.4%) roles. Table 27 shows that most participants identify as male (67.6%), are below the age of 45 ($\sim$88%), and have completed a Bachelor’s or higher degrees ($\sim$89%), with $\sim$77% in computer science, information science, software engineering, data science, and electrical, computer, privacy, or security engineering. Among the 13.1% with a business degree, $\sim$32% are in AD, $\sim$30% in AM, and $\sim$24% in RA roles. Of the 10.2% who did not have degrees within any of the above groups, $\sim$25% were in health sciences and $\sim$10% in law and psychology (among others). Most participants ($\sim$60%) have at least two years of AI development experience.
4.2 AI Perceptions and Usage
@@
To address RQ1, we examined participants’ perceptions of AI, its applications, and its impact on their work.
4.2.1 Defining AI
@@
To understand how participants perceive AI, we asked all of them to define AI within their industry. We used the NIST Taxonomy of Human-AI Use Cases [nist2024taxonomy] to categorize their responses (see Appendix B - Table 29 for more details on the taxonomy and Figure 1 for a breakdown of participants’ responses). This taxonomy does not capture some technical definitions, such as machine learning. We subsequently created a dedicated category for Machine/Deep Learning (ML/DL).
@@
Figure 1: AI Definitions Based on NIST Human-AI Taxonomy
@@
A significant number of participants in both groups (12.26% and 17.70% for Groups A and B, respectively) highlighted a wider range of use cases, which we could not map to any of the NIST categories. One participant defined AI as “It is revolutionary and can change everything in this industry.” We categorized these definitions as “Others.” Specifically, in Group B, some of these responses focus on societal or future impacts of AI (e.g., “corruptive and hugely beneficial in equal measures”). Our results indicate that those with experience in AI development (i.e., Group B) identify more technical aspects of AI and its abilities in areas such as data analysis and automation (e.g., “It is an indispensable tool for fast and efficient work that analyzes and verifies a huge volume of data. It performs work that would otherwise be impossible”).
@@
We further analyzed Group B based on their roles. “Process automation” is the most common AI definition among AM roles (41.03%), but it is considerably less common among ISec (20%) and AR (18.18%) roles. This may be because the management team is more heavily involved in process improvement, an area where AI-driven automation can have a positive impact. Participants in QA and AD roles are most likely to describe AI as “content synthesis,” which is a more technical term. QAs are also considerably more likely than other roles to mention “decision making” (27.78%). ISec roles are more than three times more likely to say “content creation” (33.33%) than other roles. This may suggest that individuals in these roles are less likely to be involved in AI development activities, and like Group A participants, see AI through the lens of contemporary narratives rather than its technical aspects. Consistent with prior work, these findings underscore the need for tools and educational frameworks that focus on privacy and security aspects of AI, as ISec team members play a critical role in addressing emerging threats that developers are often unaware of [gao2024idk, meisenbacher2024privacy].
4.2.2 AI Usage
@@
We asked both groups about the primary applications of AI in their products. As shown in Table 2, “chatbots, personal assistants, or recommender systems” are considered the most common AI applications across both groups. Group B considers “robotics and automation,” “cybersecurity,” “computer vision,” “human resources,” and “legal” more than Group A. In comparison, Group A participants are $\sim$13%, $\sim$13%, and $\sim$10% more likely to choose “translation and text generation,” “healthcare,” and “chatbots, personal assistants, or recommender systems.” These results further indicate that those without AI development experience (Group A) are more optimistic about applications tailored towards the general public, such as chatbots, while those with AI experience (Group B) consider more job-related or technical applications.
Table 2: Applications of AI in Products
@@
Group B participants could select multiple AI tools and technologies they use in their products, as shown in Figure 2. “Google Cloud AI” and “Model Provider APIs, such as OpenAI, Anthropic, LLama, etc.” are the most popular AI tools selected by the participants (49.50% and 42.52%, respectively). This suggests that development teams primarily use pre-trained models instead of building or training their own. In addition, “Azure ML,” “Pytorch,” “AWS Sagemaker,” and “Tensorflow” are more popular than “transformer-based LLMs” and other deep and machine-learning architectures.
Figure 2: Group B - Tools and Technologies Used in Product or Development
All participants were asked to discuss how often they use AI tools and what they would consider the main impact of AI tools on their work. Tables 3 and 4 show the frequency of AI usage and its impact. Most participants (i.e., 82.24% of Group A and 96.67% of Group B) use AI at least 1-3 times per week, while more than 50% of Group B uses AI tools daily. All participants consider “increased efficiency” (71.0% and 77.5% of Groups A and B) and “improved accuracy” (41.1% and 55.3% of Groups A and B) in their work. Interestingly, Group B selected “cost reduction” as the third most crucial impact of AI, with 22.2% more than Group A. In contrast, Group A identified “innovation and creativity” as the third most crucial impact. The results show that all participants recognize AI’s benefits in enhancing workplace tasks, with experienced users more frequently considering automating cost-intensive tasks.
Table 3: Frequency of AI Tool Usage
@@
Table 4: Effects of AI-Enhanced Tools on Development Process and Workflow
@@
Figure 3: Factors Limiting AI Adoption in Development Processes
Since Group A’s companies did not adopt AI in their development processes, we asked them to select factors that limit their companies’ involvement with AI. As shown in Figure 3, “privacy and security concerns” (45.19%), “lack of trust” (38.46%), and “lack of expertise”(35.58%) are the most significant factors, while “human resources” (16.35%) and “proprietary reasons” (20.19%) appear not as decisive. This highlights the importance of ethics considerations, including privacy, security, and trust, for companies that do not incorporate AI in their products. As a follow-up question, they had to describe why these factors limit their companies’ involvement. We received 106 responses, in which we identified seven common themes (see Table 5). These responses reinforced the previous finding, with $\sim$29% expressing concerns regarding security and privacy, where many reiterated worries about data sharing and consent (e.g., “customer data concerns”) or were uncertain about the security of AI systems (e.g., “if you have proprietary information and materials, you don’t want to feed it into an LLM model that will later regurgitate that information to any user”). 18.87% considered “high costs” as the main factor, while 12.26%, 11.32%, and 10.38%, mentioned a lack of “experience,” “established AI applications,” or “trust with AI systems.”
Table 5: Group A - Reasons for Limiting AI Usage in Development
@@
RQ1 - AI Perception & Usage: All participants, regardless of their experience in AI development, primarily define AI in terms of process automation or performance enhancement, report daily use of AI tools, and perceive a positive impact on efficiency and accuracy. Those with experience in AI development tend to engage with more technical aspects of AI in their work, whereas those without such experience typically utilize AI tools, such as chatbots or recommender systems, often for innovation and creativity. Concerns regarding security and privacy remain the primary reasons some companies have not yet adopted AI tools in their development.
4.3 AI Ethics Familiarity, Knowledge, and Practice
To address RQ2 and RQ3, we assessed participants’ familiarity with AI ethics principles [worsdorfer2023aibill] and governance initiatives, and how these vary across demographics.
4.3.1 Familiarity with AI Ethics Principles
We asked participants to rate their familiarity with various AI ethics principles, as shown in Figure 4. We combine the responses of somewhat, moderately, and extremely familiar. The results show that both groups are most familiar with “data protection and the right to privacy” (81.2% and 92.3%) and least familiar with “democracy and rule of law” (58.0% and 75.7%). This may be because the concept of privacy is more tangible for development teams. Those with AI development experience are more familiar with “transparency and explainability of AI systems” (89.3%) and “accountability and responsibility” (88.0%). Those in Group A are more familiar with “respect for human rights” (74.8%), while for all other principles, their familiarity is below 70%.
(a) Group A - Degree of Familiarity
(b) Group B - - Degree of Familiarity
Figure 4: Degree of Familiarity with AI Ethics Principles
@@ 4.3.1 Familiarity with AI Ethics Principles
Figure 5: Group B - Familiarity With AI Ethics Principles by Demographics
To examine whether our analysis of demographics is statistically significant, we investigated whether there is a correlation between various demographic factors, including role (H1a), location (H1b), company type (H1c), company size (H1d), and gender (H1e), and the average familiarity with AI ethics principles. As shown in Table 6, we found no evidence of a correlation between demographics and average familiarity among Group A participants. Without applying the Bonferroni correction, we observe correlations between location, company type, and gender and Group B’s average familiarity (pvalue=0.0102,0.0463p-value=0.0102,0.0463 and $0.003$), while when the Bonferroni correction is applied, only the correlation with gender remains statistically significant. Our analysis of familiarity with AI ethics principles complements prior work emphasizing the importance of diversity in demographics and expertise within AI development teams [pant2024ethics, vakkuri2022how, olson2025speaks]. Others have also highlighted the need for effective communication among development teams [stahl2022organisational], and the critical role of management in engaging their teams and stakeholders from various demographics in ethical decision-making [kawakami2024studyingup].
Table 6: P-Values for Hypotheses H1a to H1e
@@
We conducted a LASSO analysis to identify the predictive factors within demographics, perception, and practice that are most strongly associated with familiarity with AI ethics principles. For Group A, the LASSO model selected no features at the optimal regularization parameter, resulting in a null model. The model yielded a negative test R² (-0.06), indicating performance worse than a baseline model that predicts the mean. For Group B, the model indicated that increases in familiarity with AI ethics principles were most strongly associated with greater familiarity with the EU AI Act (+15.39%) (see Table 7). Additional predictors included more positive perceptions of regulatory impact (+0.1523), consideration of environmental and social responsibility in practice (+0.1403%), and the perceived effect of cost reduction for AI adoption (+0.1242%).
Figure 6: Group B - Ethics Principles in AI Development Activities
4.3.2 Practice and Experience with AI Ethics Principles
@@
We further evaluated whether Group B considers these principles in the actual AI development. Figure 6 shows the result. As expected, “data protection and the right to privacy” is considered the most important. This is probably because the concept of digital privacy has existed for much longer than other computer-ethics aspects. “Transparency and explainability of AI systems” is regarded as the second most important principle, while “democracy and rule of law” is considered the least important. The results match the degree of familiarity, as shown in Figure 7(b). We tested whether there is a correlation between the average familiarity and the use of principles in AI development (H2a). With the pvalue=0.0000p-value=0.0000, we found that there is a correlation.
Table 7: LASSO Results: Predictors of AI Ethics Principle Familiarity (Gr. B)
Predictor Coefficient Std. Error
Familiarity with Regulations
   EU AI Act $0.154^{*}$ $(0.074)$
   NIST Risk Management Framework $0.092$ $(0.106)$
   Average regulations familiarity $0.032$ $(0.129)$
Perception
   Positive perception of regulatory impact $0.152^{**}$ $(0.054)$
   Considers environment/social responsibility $0.140^{*}$ $(0.054)$
AI Practice
   AI used for cost reduction $0.124^{*}$ $(0.052)$
   Frequency of considering ethics $0.111$ $(0.062)$
   Conducts regular audits/assessments $0.101$ $(0.053)$
   Adopts AI governance frameworks $0.072$ $(0.056)$
Overall Mean 3.88
Observations 142
R² (test) 0.56
Features Selected 9 / 141
Note: $^{*}p<0.05$; $^{**}p<0.01$; $^{***}p<0.001$. Dependent variable: Average familiarity with 9 AI ethics principles (1–5 scale). Coefficients are standardized (per 1-SD change in predictor).
@@
In a more role-specific analysis, we observed that those in AM, AD, AR+AE, and RA roles consider a variety of principles in their AI development activities. For example, those in AM roles consider the principles of “respect for human rights,” “non-discrimination and freedom of privileges,” “harm prevention and beneficence,” and “fairness and justice” whereas those in AR+AE ensure that the “accountability and responsibility” is applied in the workflow. In contrast, those in QA or ISec roles do not consider these principles as much, except for “data protection and the right to privacy.” This highlights areas for improvement, as QA roles are crucial for maintaining trust in AI systems [ala-luopa2024trusting] and collaboration with ISec teams is essential to ensure the security and privacy of AI models [prybylo2024evaluating]. Participants working in academic/research institutions tend to prioritize more technical principles, such as “data protection and the right to privacy,” “fairness and justice,” and “transparency and explainability of AI systems,” when developing or researching AI models or datasets. In contrast, those employed in government agencies are likelier to emphasize abstract principles, including “respect for human rights,” “non-discrimination and freedom of privileges,” and “harm prevention and beneficence.” Those in EU+EEA+UK consider “data protection and the right to privacy” and “democracy and rule of law” more than those who work in North America and other locations. This is expected since privacy is considered a fundamental (human) right and is regulated by the General Data Protection Regulation (GDPR) [GDPR]. In line with the above results, female participants consider AI ethics principles in their AI development activities more than their male counterparts.
@@
We asked Group B how often they consider AI ethics principles in their work. Most participants consider them “Always” (23.92%) or “Often” (44.52%), while only 9.97% and 2.33%, respectively, consider them “Rarely” or “Never.” Our results show that those in QA (92.3%), ISec (91.7%), AM (91.5%), and AR+AE (90%) roles are more likely to consider a subset of AI ethics principles in their work at least occasionally, whereas those in AD (85.4%) and RA (75%) roles report slightly lower engagement. Those with a BSc. or higher degree are $\sim$10% more likely to consider AI ethics principles in their work than those without. Similar to prior observations, female participants include AI ethics principles in their work $\sim$13% more than their male counterparts.
We conducted a separate LASSO analysis to examine the predictive factors most strongly associated with how frequently AI developers consider ethics principles in their daily work (ranging from never to always). The results show that more frequent ethical consideration is most strongly associated with higher overall familiarity with AI ethics principles (+0.281), more positive perceptions of regulatory impact (+0.226), and familiarity with the US Executive Order on Safe AI (+0.210) (see Table 8).
Table 8: LASSO Results: Predictors of Ethics Consideration Frequency (Gr. B)
Predictor Coefficient Std. Error
Familiarity with AI Ethics Principles
   Average principle familiarity $0.281^{**}$ $(0.088)$
Perception
   Positive perception of regulatory impact $0.226^{**}$ $(0.076)$
Familiarity with Regulations
   US Executive Order on Safe AI $0.210^{*}$ $(0.083)$
Overall Mean 3.73
Observations 137
R² (test) 0.40
Features Selected 3 / 150
Note: $^{*}p<0.05$; $^{**}p<0.01$; $^{***}p<0.001$. Dependent variable: Frequency of considering AI ethics principles in development (1–5 scale, from never to always). Coefficients are standardized (per 1-SD change in predictor).
@@
RQ2 & RQ3 - AI Ethics Principles: Our results highlight that the degree of familiarity with AI ethics principles - and the extent to which they are considered in practice - varies across demographic factors. AM roles demonstrate greater familiarity with most principles and are more likely to incorporate them into their work. In contrast, those in QA or ISec roles tend to be the least familiar with them, with the exception of “data protection and the right to privacy.” Both location and company type influence levels of familiarity and application. Overall, female participants not only reported higher familiarity with all principles but also indicated more consistent integration of these principles into their workflows. Our LASSO analysis reveals that familiarity with AI governance initiatives, particularly the EU AI Act, is the most significant predictive factor of familiarity with AI ethics principles.
4.3.3 Familiarity with AI Governance Initiatives